# Factor Models of Matrix-Valued Time Series: Nonstationarity and Cointegration

### v1

@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 introduces the matrix factor model with discussion on model identifiability, and the mPCA and mPANIC estimation methods. Section 3 presents the convergence properties of the developed estimators with assumptions and remarks. They are extended to the case of heterogeneous factor strength in Section 4. Section 5 reports both the simulation and empirical studies. Section 6 concludes the paper. Appendix A contains the proofs of the main asymptotic results and Appendix B contains some technical lemmas with proofs. Throughout the paper, we let ${\sf vec}(\cdot)$ be the vectorization of a matrix and ${\sf vec}^{-1}(\cdot)$ inverse of the vectorization operator. We use $\otimes$ and $\odot$ to denote the Kronecker and Hadamard products, respectively. Let $\|\cdot\|$ be the Euclidean norm of a vector or operator norm of a matrix, and $\|\cdot\|_{F}$ the matrix Frobenius norm. Let $\psi_{k}(\mathbf{A})$ denote the $k$-th largest eigenvalue of a positive semidefinite matrix $\mathbf{A}$. For notational simplicity, we write $\max\{a,b\}$” as $a\vee b$”, $\min\{a,b\}$” as $a\wedge b$”, and “with probability approaching one” as “w.p.a.1”.
@@ 2.1 Model formulation
We are interested in the following two scenarios for latent factors and idiosyncratic components: (i) the nonstationary matrix factors are of full rank and the idiosyncratic components are stationary over time, and (ii) the matrix factors are cointegrated and the idiosyncratic components are either stationary or nonstationary. The nonstationarity of matrix time series in scenario (i) is fully captured by the common stochastic trends contained in matrix factors, whereas appropriate rotation of the cointegrated factors in scenario (ii) can separate out stationary and nonstationary factors. In the following two subsections, we provide detailed model assumptions and propose two types of PCA-based estimation techniques for these two scenarios, respectively.
@@
where $\widetilde{\lambda}_{R,1}$ is the maximum eigenvalue of $T^{-1}\widehat{\boldsymbol{\Omega}}_{R}$, $\widetilde{\boldsymbol{V}}_{R}$ is an $r_{1}\times r_{1}$ diagonal matrix with the diagonal entries being the $r_{1}$ largest eigenvalues of $T^{-1}\widehat{\boldsymbol{\Omega}}_{R}$ (in a descending order) whereas $\widetilde{\boldsymbol{V}}_{C}$ is an $r_{2}\times r_{2}$ diagonal matrix with the diagonal entries being the $r_{2}$ largest eigenvalues of $T^{-1}\widehat{\boldsymbol{\Omega}}_{C}$. The adaptive estimation formula in (2.4) is due to possible existence of weak factors with the extra adaptive factors $\widetilde{\lambda}_{R,1}^{1/2}$, $\widetilde{\boldsymbol{V}}_{R}^{-1/2}$ and $\widetilde{\boldsymbol{V}}_{C}^{-1/2}$ related to the factor strength, see Assumptions 3.1(ii) and 4.1. When both the row and column factors are assumed to be strong, the estimation formula in (2.4) can be simplified to $\widehat{\boldsymbol{F}}_{t}=(p_{1}p_{2})^{-1/2}\widehat{\boldsymbol{R}}^{{}^{\intercal}}{\boldsymbol{X}}_{t}\widehat{\boldsymbol{C}}$.
@@ 3 Large-sample theory
In this section, we give some regularity conditions and establish the convergence properties for the mPCA and mPANIC estimators developed in Section 2. The factor strength is assumed to be homogeneous over (weak) factors in this section and the extension of theory to the case of heterogeneous factor strength will be explored in Section 4.
@@ 3.1 Convergence theory for mPCA estimation
(ii) Letting $\rho_{R,t,(i_{1},i_{2})}=\frac{1}{p_{2}}\sum_{j=1}^{p_{2}}{\sf E}[e_{t,(i_{1},j)}e_{t,(i_{2},j)}]$, $\rho_{C,t,(j_{1},j_{2})}=\frac{1}{p_{1}}\sum_{i=1}^{p_{1}}{\sf E}[e_{t,(i,j_{1})}e_{t,(i,j_{2})}]$
@@
where $\widetilde{\mathbf{u}}_{t}=\sum_{j=0}^{\infty}\widetilde{\boldsymbol{A}}_{j}{\boldsymbol{\eta}}_{t-j}$ with $\widetilde{\boldsymbol{A}}_{j}=\sum_{k=j+1}^{\infty}{\boldsymbol{A}}_{k}$. By (3.1), $\{\widetilde{\mathbf{u}}_{t}\}$ is a stationary vector linear process and $\overline{\boldsymbol{A}}{\boldsymbol{\Sigma}}_{\eta}\overline{\boldsymbol{A}}^{{}^{\intercal}}$ is of full rank. Hence, ${\boldsymbol{F}}_{t}$ is a matrix of full-rank integrated random elements. In fact, with (3.5) and Gaussian approximation theorem in Theorem 4 of Götze and Zaitsev (2009), we may show that
@@
More details are provided in the proof of Lemma B.3. Assumption 3.1(ii) allows the existence of weak factors with $\alpha_{R}$ and $\alpha_{C}$ representing the strength of row and column factors, respectively (e.g., Lam et al, 2011; Chen et al, 2019; Wang et al, 2019). When $\alpha_{R}=\alpha_{C}=1$, the row and column factors have impacts on the majority of matrix time series entries, and are thus called “strong factors” (e.g., Chen and Fan, 2023). In addition, (3.2) and (3.3) indicate that the nonstationary factors have homogeneous strength. We will consider the extension to the setting with heterogeneous factor strength in Section 4 below.
(ii) Assumption 3.2 contains some high-level conditions which are mild restrictions on temporal and cross-row (or column) correlation. Similar assumptions can also be found in Chen and Fan (2023) and they can be seen as a matrix extension of those conditions used by Bai and Ng (2002) and Bai (2004). The restriction on temporal dependence can be justified if we impose a mixing dependence condition on $\{{\boldsymbol{E}}_{t}\}$. From Assumption 3.2, the idiosyncratic components are not necessarily stationary and are allowed to be heteroskedatic over time. Assumption 3.2(i) indicates that $\{{\boldsymbol{E}}_{t}\}$ and $\{{\boldsymbol{F}}_{t}\}$ are mutually independent, but this independence restriction can be relaxed at the cost of more lengthy proofs.
@@
where $\widehat{\boldsymbol{V}}_{R}$ is an $r_{1}\times r_{1}$ diagonal matrix with the diagonal entries being the $r_{1}$ largest eigenvalues of $(p_{1}^{\alpha_{R}}p_{2}^{\alpha_{C}}T)^{-1}\widehat{\boldsymbol{\Omega}}_{R}$ (in a descending order) and $\widehat{\boldsymbol{V}}_{C}$ is an $r_{2}\times r_{2}$ diagonal matrix with the diagonal entries being the $r_{2}$ largest eigenvalues of $(p_{1}^{\alpha_{R}}p_{2}^{\alpha_{C}}T)^{-1}\widehat{\boldsymbol{\Omega}}_{C}$. Lemma B.5 in Appendix B shows that both $\widehat{\boldsymbol{H}}_{R}$ and $\widehat{\boldsymbol{H}}_{C}$ are invertible w.p.a.1, and derives their respective limits. The following theorem gives the (normalized) mean squared convergence rates for the mPCA estimates: $\widehat{\boldsymbol{R}}$ and $\widehat{\boldsymbol{C}}$.
@@
The rate $\varphi_{1}(p_{1},p_{2},T)$ is due to the uniform convergence for the first term on RHS of (3.10), making use of Theorem 3.1, whereas the rate $\varphi_{2}(p_{1},p_{2},T)$ is due to the uniform convergence of the second term on RHS of (3.10) based on a combination of Theorem 3.1 and Lemma B.6 in Appendix B.
@@
(ii) Unlike the stationary case (e.g., Freyaldenhoven, 2022), consistent estimation of the number of factors in our framework does not require the minimum factor strengths $\alpha_{R,r_{1}}$ and $\alpha_{C,r_{2}}$ to exceed a fixed threshold, as long as Assumption 3.3(i) holds. Intuitively, this is due to the fact that our method directly targets the nonstationary matrix time series and strong signals in the common stochastic trends dominate those in the matrix idiosyncratic components, leading to much faster convergence as illustrated in Theorem 3.1 and Remark 3.2.
@@ 3.2 Convergence theory for mPANIC estimation
We next derive the convergence properties for the mPANIC estimator developed in Section 2.3, and start with some regularity conditions.
@@
Assumption 3.5(i) is crucial to derive the classic Granger’s representation for latent factors (e.g., Johansen, 1995; Cheng and Phillips, 2009; Li and Xiao, 2024), see (B.20) in Appendix B. The independence condition on $\{{\boldsymbol{V}}_{t}\}$ in Assumption 3.5(ii) may be replaced by stationary martingale differences. Assumption 3.5(iii) is comparable to Assumption 3.1(iii) and a similar condition has been commonly used in the literature on factor model estimation theory. Assumption 3.6 is analogous to Assumption 3.2, containing some high-level restrictions on temporal, cross-row and cross-column dependence. Assumption 3.7 implies that $\alpha_{R}$ and $\alpha_{C}$ must be larger than $1/2$, which is more restrictive that that in Section 3.1.
@@
where $\check{\boldsymbol{V}}_{R}$ is an $r_{1}\times r_{1}$ diagonal matrix with the diagonal entries being the $r_{1}$ largest eigenvalues of $(p_{1}^{\alpha_{R}}p_{2}^{\alpha_{C}})^{-1}\overline{\boldsymbol{\Omega}}_{R}$ (in a descending order), and $\check{\boldsymbol{V}}_{C}$ is defined similarly with $r_{1}$ and $\overline{\boldsymbol{\Omega}}_{R}$ replaced by $r_{2}$ and $\overline{\boldsymbol{\Omega}}_{C}$, respectively. The following theorem gives the (normalized) mean squared convergence rates for mPANIC estimates of row and column factor loadings.
@@
Suppose that Assumptions 3.1(ii), 3.53.7 are satisfied. Then we have
@@
Assumption 3.8(i) restricts the relationship between the dimensions $p_{1}$, $p_{2}$ and $T$, indicating that $s_{R}\ll p_{1}^{\alpha_{R}}$ and $s_{C}\ll p_{2}^{\alpha_{C}}$. When $\alpha_{R}=\alpha_{C}=1$, Assumption 3.8(i) can be simplified to
@@
Suppose that Assumptions 3.1(ii), 3.53.8 are satisfied. The following uniform convergence property holds:
@@ 4.1 Convergence of mPCA with heterogeneous factor strength
Assumptions 4.1 and 4.2 are similar in spirit to Assumptions 3.1(ii)(iii), 3.2(iii) in Section 3.1. In particular, Assumption 4.1(i) characterizes the heterogeneous factor strengths via the diagonal matrices $\mathbf{B}_{R}$ and $\mathbf{B}_{C}$. Assuming ${\boldsymbol{\Sigma}}_{R}^{*}$ and ${\boldsymbol{\Sigma}}_{C}^{*}$ to be diagonal allows for the identification of factors with different strength orders and facilitates the choice of suitable normalizers, ensuring that the sample eigenvalue matrices $\widehat{\boldsymbol{V}}_{R}^{\ast}$ and $\widehat{\boldsymbol{V}}_{C}^{\ast}$ (to be defined below) are positive definite. This condition is also adopted by some recent works on vector and matrix factor models with weak loadings (or factors), e.g., Assumption A2’(ii) in Bai and Ng (2023) and Assumption (V1) in Chen and Lam (2024b). As pointed out by Chen and Lam (2024b), this requirement is not restrictive, since weak factor models with heterogeneous strengths and non-diagonal limiting matrices are observationally equivalent to those with diagonal ones. Assumption 4.3 requires a more stringent requirement on the relative growth rates of $p_{1}$, $p_{2}$, and $T$. In particular, the precise condition depends on the degree of heterogeneity between the strongest and weakest factor strengths, similar to Assumption A4 in Bai and Ng (2023) for vectorized factor models.
Due to the heterogeneous factor strength, a new normalization matrix is required to properly scale the sample moments of row/column factor loadings, see Assumption 4.1(i). Accordingly, the rotation matrices also need to be modified in an appropriate manner, i.e.,
@@
where $\widehat{\boldsymbol{V}}_{R}^{\ast}=p_{2}^{-\alpha_{C,1}}\mathbf{B}_{R}^{-2}\widetilde{\boldsymbol{V}}_{R}$, $\widehat{\boldsymbol{V}}_{C}^{\ast}=p_{1}^{-\alpha_{R,1}}\mathbf{B}_{C}^{-2}\widetilde{\boldsymbol{V}}_{C}$, $\widetilde{\boldsymbol{V}}_{R}$ and $\widetilde{\boldsymbol{V}}_{C}$ are defined as in (2.4). Lemma B.15 in Appendix B shows that both $\widehat{\boldsymbol{H}}_{R}$ and $\widehat{\boldsymbol{H}}_{C}$ are invertible w.p.a.1. The following theorem extends Theorems 3.13.3 to the case of varying factor strength.

### v2

@@ 1 Introduction
The rest of the paper is organized as follows. Section 2 introduces the matrix factor model with discussion on model identifiability, and the mPCA and mPANIC estimation methods. Section 3 presents the convergence properties of the developed estimators with assumptions and remarks. They are extended to the case of heterogeneous factor strength in Section 4. Section 5 reports both the simulation and empirical studies. Section 6 concludes the paper. The supplemental document contains the proofs of the main asymptotic results and some technical lemmas with proofs. Throughout the paper, we let ${\sf vec}(\cdot)$ be the vectorization of a matrix and ${\sf vec}^{-1}(\cdot)$ inverse of the vectorization operator. We use $\otimes$ and $\odot$ to denote the Kronecker and Hadamard products, respectively. Let $\|\cdot\|$ be the Euclidean norm of a vector or operator norm of a matrix, and $\|\cdot\|_{F}$ the matrix Frobenius norm. Let $\psi_{k}(\mathbf{A})$ denote the $k$-th largest eigenvalue of a positive semidefinite matrix $\mathbf{A}$. For notational simplicity, we write $\max\{a,b\}$” as $a\vee b$”, $\min\{a,b\}$” as $a\wedge b$”, and “with probability approaching one” as “w.p.a.1”.
@@ 2.1 Model formulation
We are interested in the following two scenarios for latent factors and idiosyncratic components: (i) the nonstationary matrix factors are of full rank and the idiosyncratic components are stationary over time, and (ii) the matrix factors are cointegrated and the idiosyncratic components are either stationary or nonstationary. The nonstationarity of matrix time series in scenario (i) is fully captured by the common stochastic trends contained in matrix factors, whereas appropriate rotation of the cointegrated factors in scenario (ii) can separate out stationary and nonstationary factors. In the following two subsections, we provide detailed model settings and propose two types of PCA-based estimation techniques for these two scenarios, respectively.
@@
where $\widetilde{\lambda}_{R,1}$ is the maximum eigenvalue of $T^{-1}\widehat{\boldsymbol{\Omega}}_{R}$, $\widetilde{\boldsymbol{V}}_{R}$ is an $r_{1}\times r_{1}$ diagonal matrix with the diagonal entries being the $r_{1}$ largest eigenvalues of $T^{-1}\widehat{\boldsymbol{\Omega}}_{R}$ (in a descending order) whereas $\widetilde{\boldsymbol{V}}_{C}$ is an $r_{2}\times r_{2}$ diagonal matrix with the diagonal entries being the $r_{2}$ largest eigenvalues of $T^{-1}\widehat{\boldsymbol{\Omega}}_{C}$. The adaptive estimation formula in (2.4) is due to possible existence of weak factors with the extra adaptive factors $\widetilde{\lambda}_{R,1}^{1/2}$, $\widetilde{\boldsymbol{V}}_{R}^{-1/2}$ and $\widetilde{\boldsymbol{V}}_{C}^{-1/2}$ related to the (varying) factor strength, see Assumptions 3.1(ii) and 4.1(i). When both the row and column factors are assumed to be strong, the estimation formula in (2.4) can be simplified to $\widehat{\boldsymbol{F}}_{t}=(p_{1}p_{2})^{-1/2}\widehat{\boldsymbol{R}}^{{}^{\intercal}}{\boldsymbol{X}}_{t}\widehat{\boldsymbol{C}}$.
@@ 3 Large-sample theory
In this section, we give some regularity conditions and establish the convergence properties for the mPCA and mPANIC estimators developed in Section 2. The factor strength is assumed to be homogeneous over (weak) factors in this section and the extension to the case of heterogeneous factor strength will be explored in Section 4.
@@ 3.1 Convergence theory for mPCA estimation
(ii) Letting $\rho_{R,t,(i_{1},i_{2})}=\frac{1}{p_{2}}\sum_{j=1}^{p_{2}}{\sf E}[e_{t,(i_{1},j)}e_{t,(i_{2},j)}]$ and $\rho_{C,t,(j_{1},j_{2})}=\frac{1}{p_{1}}\sum_{i=1}^{p_{1}}{\sf E}[e_{t,(i,j_{1})}e_{t,(i,j_{2})}]$,
@@
where $\widetilde{\mathbf{u}}_{t}=\sum_{j=0}^{\infty}\widetilde{\boldsymbol{A}}_{j}{\boldsymbol{\eta}}_{t-j}$ with $\widetilde{\boldsymbol{A}}_{j}=\sum_{k=j+1}^{\infty}{\boldsymbol{A}}_{k}$. By (3.1), $\{\widetilde{\mathbf{u}}_{t}\}$ is a stationary vector linear process and $\overline{\boldsymbol{A}}{\boldsymbol{\Sigma}}_{\eta}\overline{\boldsymbol{A}}^{{}^{\intercal}}$ is of full rank. Hence, ${\boldsymbol{F}}_{t}$ is a matrix of full-rank integrated random elements. In fact, with (3.5) and Gaussian approximation theorem in Götze and Zaitsev (2009), we may show that
@@
More details are provided in the proof of Lemma B.3 available in the supplement. Assumption 3.1(ii) allows the existence of weak factors with $\alpha_{R}$ and $\alpha_{C}$ representing the strength of row and column factors, respectively (e.g., Lam et al, 2011; Chen et al, 2019; Wang et al, 2019). When $\alpha_{R}=\alpha_{C}=1$, the row and column factors have impacts on the majority of matrix time series entries, and are thus called “strong factors” (e.g., Chen and Fan, 2023). In addition, (3.2) and (3.3) indicate that the nonstationary factors have homogeneous strength. We will consider the extension to the setting with heterogeneous factor strength in Section 4 below.
(ii) Assumption 3.2 contains some high-level conditions which are mild restrictions on temporal and cross-row (or column) correlation. Similar assumptions can also be found in Chen and Fan (2023) and they can be seen as a matrix extension of those conditions used by Bai and Ng (2002) and Bai (2004). The restriction on temporal dependence can be justified if we impose a mixing dependence condition on $\{{\boldsymbol{E}}_{t}\}$. From Assumption 3.2, the idiosyncratic components are allowed to be heteroskedatic over time. Assumption 3.2(i) indicates that $\{{\boldsymbol{E}}_{t}\}$ and $\{{\boldsymbol{F}}_{t}\}$ are mutually independent, but this independence restriction can be relaxed at the cost of more lengthy proofs.
@@
where $\widehat{\boldsymbol{V}}_{R}$ is an $r_{1}\times r_{1}$ diagonal matrix with the diagonal entries being the $r_{1}$ largest eigenvalues of $(p_{1}^{\alpha_{R}}p_{2}^{\alpha_{C}}T)^{-1}\widehat{\boldsymbol{\Omega}}_{R}$ (in a descending order) and $\widehat{\boldsymbol{V}}_{C}$ is an $r_{2}\times r_{2}$ diagonal matrix with the diagonal entries being the $r_{2}$ largest eigenvalues of $(p_{1}^{\alpha_{R}}p_{2}^{\alpha_{C}}T)^{-1}\widehat{\boldsymbol{\Omega}}_{C}$. Lemma B.5 in Appendix B of the supplement shows that both $\widehat{\boldsymbol{H}}_{R}$ and $\widehat{\boldsymbol{H}}_{C}$ are invertible w.p.a.1, and derives their respective limits. The following theorem gives the mean squared convergence rates for the (normalized) mPCA estimates.
@@
The rate $\varphi_{1}(p_{1},p_{2},T)$ is due to the uniform convergence for the first term on RHS of (3.10), making use of Theorem 3.1, whereas the rate $\varphi_{2}(p_{1},p_{2},T)$ is due to the uniform convergence of the second term on RHS of (3.10) based on a combination of Theorem 3.1 and Lemma B.6 in Appendix B of the supplement.
@@
(ii) Unlike the stationary case (e.g., Freyaldenhoven, 2022), consistent estimation of the number of factors in our framework does not require the factor strengths $\alpha_{R}$ and $\alpha_{C}$ to exceed a fixed threshold, as long as Assumption 3.3(i) holds. Intuitively, this is due to the fact that our method directly targets the nonstationary matrix time series and strong signals in the common stochastic trends dominate those in the matrix idiosyncratic components, leading to much faster convergence as illustrated in Theorem 3.1 and Remark 3.2.
@@ 3.2 Convergence theory for mPANIC estimation
We next derive the convergence properties for the mPANIC estimator developed in Section 2.3.
@@
Assumption 3.5(i) is crucial to derive the classic Granger’s representation for latent factors (e.g., Johansen, 1995; Cheng and Phillips, 2009; Li and Xiao, 2024), see (B.20) in the supplement. The independence condition on $\{{\boldsymbol{V}}_{t}\}$ in Assumption 3.5(ii) may be replaced by stationary martingale differences. Assumption 3.5(iii) is comparable to Assumption 3.1(iii) and a similar condition has been commonly used in the literature on factor model estimation theory. Assumption 3.6 is analogous to Assumption 3.2, containing some high-level restrictions on temporal, cross-row and cross-column dependence. Assumption 3.7 implies that $\alpha_{R}$ and $\alpha_{C}$ must be larger than $1/2$, which is more restrictive that that in Section 3.1.
@@
where $\check{\boldsymbol{V}}_{R}$ is an $r_{1}\times r_{1}$ diagonal matrix with the diagonal entries being the $r_{1}$ largest eigenvalues of $(p_{1}^{\alpha_{R}}p_{2}^{\alpha_{C}})^{-1}\overline{\boldsymbol{\Omega}}_{R}$ (in a descending order), and $\check{\boldsymbol{V}}_{C}$ is defined similarly with $r_{1}$ and $\overline{\boldsymbol{\Omega}}_{R}$ replaced by $r_{2}$ and $\overline{\boldsymbol{\Omega}}_{C}$, respectively. The following theorem gives the mean squared convergence rates for (normalized) mPANIC estimates of row and column factor loadings.
@@
Suppose that Assumptions 3.1(ii) and 3.53.7 are satisfied. Then we have
@@
Assumption 3.8(i) restricts the relationship between the dimensions $p_{1}$, $p_{2}$ and $T$, indicating that $s_{R}\ll p_{1}^{\alpha_{R}}$ and $s_{C}\ll p_{2}^{\alpha_{C}}$. When $\alpha_{R}=\alpha_{C}=1$, Assumption 3.8(i) is simplified to
@@
Suppose that Assumptions 3.1(ii) and 3.53.8 are satisfied. The following uniform convergence property holds:
@@ 4.1 Convergence of mPCA with heterogeneous factor strength
Assumptions 4.1 and 4.2 are similar in spirit to Assumptions 3.1(ii)(iii) and 3.2(iii) in Section 3.1. In particular, Assumption 4.1(i) characterizes heterogeneous factor strengths via the diagonal matrices $\mathbf{B}_{R}$ and $\mathbf{B}_{C}$. Assuming ${\boldsymbol{\Sigma}}_{R}^{*}$ and ${\boldsymbol{\Sigma}}_{C}^{*}$ to be diagonal allows for the identification of factors with different strength orders and facilitates the choice of suitable normalizers, ensuring that the sample eigenvalue matrices $\widehat{\boldsymbol{V}}_{R}^{\ast}$ and $\widehat{\boldsymbol{V}}_{C}^{\ast}$ (to be defined below) are positive definite. This condition is also adopted by some recent works on vector and matrix factor models with weak loadings (or factors), e.g., Assumption A2’(ii) in Bai and Ng (2023) and Assumption (V1) in Chen and Lam (2024b). As pointed out by Chen and Lam (2024b), this requirement is not restrictive, since weak factor models with heterogeneous strengths and non-diagonal limiting matrices are observationally equivalent to those with diagonal ones. Assumption 4.3 requires a more stringent requirement on the relative growth rates of $p_{1}$, $p_{2}$, and $T$. In particular, the precise condition also depends on the degree of heterogeneity between the strongest and weakest factor strengths, similar to Assumption A4 in Bai and Ng (2023) for vectorized factor models.
Due to the heterogeneous factor strength, a new normalization matrix is required to properly scale the sample moments of row/column factor loadings, see Assumption 4.1(i). Accordingly, the rotation matrices also need to be modified, i.e.,
@@
where $\widehat{\boldsymbol{V}}_{R}^{\ast}=p_{2}^{-\alpha_{C,1}}\mathbf{B}_{R}^{-2}\widetilde{\boldsymbol{V}}_{R}$, $\widehat{\boldsymbol{V}}_{C}^{\ast}=p_{1}^{-\alpha_{R,1}}\mathbf{B}_{C}^{-2}\widetilde{\boldsymbol{V}}_{C}$, $\widetilde{\boldsymbol{V}}_{R}$ and $\widetilde{\boldsymbol{V}}_{C}$ are defined as in (2.4). Lemma B.15 in Appendix B of the supplement shows that both $\widehat{\boldsymbol{H}}_{R}$ and $\widehat{\boldsymbol{H}}_{C}$ are invertible w.p.a.1. The following theorem extends Theorems 3.13.3 to the case of varying factor strength.